# Preserving Privacy Without Compromising Accuracy: Machine Unlearning for Handwritten Text Recognition

### v1

@@ 4.6.1 Membership Inference Evaluation
5 Conclusion and Future Directions
In this study, we introduce a novel two-stage machine unlearning approach tailored for multi-head transformer-based handwriting text recognition (HTR) models. Our method effectively mitigates the retention of user-specific information while preserving recognition performance. The first stage leverages neural pruning to systematically eliminate dominant information associated with the target data by identifying and removing neurons with high importance scores, determined through the activation ratio between the forget set and the retain set. This targeted pruning minimizes the impact of the forget set on model predictions. In the second stage, we apply random labeling through the writer classification head, ensuring that user-identifiable patterns are effectively erased without compromising overall recognition accuracy.
To assess the robustness of our approach, we conduct extensive membership inference attack evaluations, demonstrating its efficiency in protecting user privacy with a minimal number of unlearning iterations. Our findings highlight the effectiveness of our method in reducing memorization while maintaining model usability, making it a practical solution for real-world applications. As the first comprehensive exploration of machine unlearning in the context of HTR, this research significantly contributes to the document analysis community by bridging the gap between privacy-preserving methodologies and handwriting recognition. Our framework not only advances privacy-aware document understanding but also sets a foundation for future developments in secure and adaptive HTR systems.
Looking ahead, this work opens several promising avenues for future research. A key direction involves the development of layout-aware machine unlearning techniques that move beyond isolated text regions to encompass structured document understanding. Such approaches should aim to selectively remove sensitive content while preserving the visual and semantic integrity of complex layouts including tables, figures, and spatial hierarchies.
Another compelling research trajectory is multilingual machine unlearning, which focuses on selectively forgetting specific languages in multilingual HTR systems. This includes investigating potential cross-linguistic interference and understanding how unlearning one language may impact the model’s performance on others, especially in low-resource or script-diverse settings.
Finally, a broader and more forward-looking direction is the extension to downstream applications such as document visual question answering (DocVQA) [35]. In this context, integrating the proposed prune-unlearn mechanisms into retrieval-based DocVQA systems [36] could enable models to be updated in response to user requests, such that they can no longer retrieve certain documents. This would effectively prevent those documents from being accessed or used in future question-answering tasks, thereby providing stronger end-to-end privacy guarantees.
@@ Acknowledgements
This work has been supported by the Beatriu de Pinós del Departament de Recerca i Universitats de la Generalitat de Catalunya (2022 BP 00256), the predoctoral program AGAUR-FI ajuts (2024 FI-3 00065) Joan Oró, which is backed by the Secretariat of Universities and Research of the Department of Research and Universities of the Generalitat of Catalonia, as well as the European Social Plus Fund, European Lighthouse on Safe and Secure AI (ELSA) from the European Union’s Horizon Europe programme under grant agreement No 101070617, Ramon y Cajal research fellowship RYC2020-030777-I / AEI / 10.13039/501100011033, Riksbankens Jubileumsfond, grant M24-0028: Echoes of History: Analysis and Decipherment of Historical Writings (DESCRYPT), the Spanish projects CNS2022-135947 (DOLORES, 2023-2025) and PID2021-126808OB-I00 (GRAIL, 2022-2024), the Consolidated Research Group 2021 SGR 01559 from the Research and University Department of the Catalan Government, and PID2023-146426NB-100 funded by MCIU/AEI/10.13039/501100011033 and FSE+.

### v2

@@ 4.6.1 Random Labeling Experiments
Thus, we can now address research question [RQ3] by concluding that using machine unlearning techniques in the writer classification head effectively removes user-identifiable information without negatively impacting recognition performance.
4.6.2 Fisher Forgetting Experiments
Fisher Forgetting is applied once over the entire forget set to compute both the gradient $g_{f}$ and the diagonal Fisher matrix. Since the forget set contains 2,007 samples and the batch size is 64, this step is equivalent to approximately 32 iterations. Given the total budget of 10,000 iterations, we use 32 iterations for Fisher Forgetting and allocate the remaining 9,968 iterations for fine-tuning.
Table 12: Experiments with Fisher Forgetting applied to both the baseline and pruned models.
Method Iter. Sparsity Forget Set Retain Set Test Set
ACC CER WER ACC CER WER CER WER
Baseline $M$ 0 0% 100.00 1.80 1.89 100.00 1.29 1.75 13.23 34.90
+Fisher 32 0% 5.38 52.45 76.63 6.02 50.60 75.95 60.63 81.41
+FT 9,968 0% 7.67 11.73 8.72 99.98 8.64 8.40 25.08 47.65
Pruned $M^{*}$ 0 17.45% 20.03 58.29 84.16 35.50 60.44 85.26 66.81 88.32
+Fisher 32 8.28% 6.08 584.07 100.00 6.37 581.22 100.00 593.86 100.00
+FT 9,968 7.39% 8.87 5.21 9.77 99.99 2.52 3.82 15.16 37.38
From Table 12, we observe that applying Fisher Forgetting alone leads to a performance drop in accuracy, CER, and WER for both the forget set and the retain set. This degradation is likely due to the highly entangled nature of information in sequence recognition models, which makes targeted forgetting more challenging compared to simpler classification tasks.
To recover overall performance after unlearning, the model is further fine-tuned on the retain set. Surprisingly, despite the forget set being excluded during this fine-tuning phase, the writer classification accuracy on the forget set slightly improves. This indicates that Fisher Forgetting alone may not have fully removed the forget set’s influence. One possible explanation is that latent representations associated with the forget set remain in the model and are unintentionally reinforced during retain-set fine-tuning.
Moreover, the MIA success rate on the forget set is 64%, which is above the random guessing baseline of 50%. This indicates a failure to fully achieve the unlearning objective from a privacy perspective, as the model still partially treats the forget set as training members. While Fisher Forgetting may reduce some direct memorization, residual user-identifiable signals persist, highlighting its limitations in fully erasing sensitive information.
4.6.3 Amnesiac Unlearning Experiments
Amnesiac Unlearning [33] reverses the influence of the forget set by applying negative gradient updates scaled by a factor $\alpha$, effectively approximating the inverse of the original training dynamics. The scaling factor $\alpha$ is set to $1\times 10^{-6}$.
Table 13: Experiments with Amnesiac Unlearning applied to both the baseline and pruned models.
Method Iter. Sparsity Forget Set Retain Set Test Set
ACC CER WER ACC CER WER CER WER
Baseline $M$ 0 0% 100.00 1.80 1.89 100.00 1.29 1.75 13.23 34.90
+Amnesiac 32 0% 3.14 69.28 89.24 0.10 68.35 88.52 76.05 91.60
+FT 9,968 0% 4.79 7.48 7.42 99.96 4.92 5.56 20.45 45.18
Pruned $M^{*}$ 0 17.45% 20.03 58.29 84.16 35.50 60.44 85.26 66.81 88.32
+Amnesiac 32 7.84% 0.00 88.38 98.75 0.39 89.18 98.71 89.44 98.57
+FT 9,968 7.40% 7.47 4.88 8.87 99.99 3.92 4.87 16.53 38.77
From Table 13, we observe that applying Amnesiac Unlearning alone leads to a drop in accuracy, CER, and WER for both the forget set and the retain set. This performance degradation highlights the highly entangled nature of learned representations between the forget and retain sets in handwriting recognition tasks. The difficulty of isolating user-specific information without affecting shared representations makes effective unlearning particularly challenging in this context.
Interestingly, we find that the pruning stage significantly strengthens the impact of Amnesiac Unlearning. Specifically, pruning leads to a strong removal of information, reducing the writer classification accuracy on the forget set to 0% at this point. However, after fine-tuning the model on the retain set, the accuracy on the forget set increases to 7.47%. This recovery suggests that, although pruning initially removes the forget set information, some hidden traces may still remain in the model and get activated again during fine-tuning. This shows that Amnesiac Unlearning, even when combined with pruning, may still struggle to fully prevent the return of user-identifiable patterns, likely due to the shared features between the forget and retain sets.
Supporting this observation, the MIA success rate on the forget set reaches 67%, which is significantly above the random baseline of 50%. This result indicates that the model continues to treat the forget set as part of its training data, confirming that the unlearning process is incomplete from a privacy standpoint.
4.6.4 DELETE Experiments
We further evaluate the effectiveness of the DELETE method [34] in our prune-unlearn pipeline. To ensure a fair comparison with other machine unlearning techniques, we follow the same experimental setup by limiting the total training iterations to 10,000. Specifically, we allocate 5,000 iterations to the DELETE unlearning phase, during which the model is optimized to remove the influence of the forget set. The remaining 5,000 iterations are dedicated to fine-tuning the model exclusively on the retain set to recover any potential loss in task performance.
Table 14: Experiments with DELETE applied to both the baseline and pruned models.
Method Iter. Sparsity Forget Set Retain Set Test Set
ACC CER WER ACC CER WER CER WER
Baseline $M$ 0 0% 100.00 1.80 1.89 100.00 1.29 1.75 13.23 34.90
+DELETE 5,000 0% 0.40 21.70 41.16 99.93 16.84 26.48 36.05 65.03
+FT 5,000 0% 0.00 7.54 7.67 99.99 7.16 7.25 22.11 46.06
Pruned $M^{*}$ 0 17.45% 20.03 58.29 84.16 35.50 60.44 85.26 66.81 88.32
+DELETE 5,000 7.48% 0.00 62.26 87.54 92.85 61.57 87.82 70.24 90.89
+FT 5,000 7.43% 0.00 7.01 11.16 99.99 4.70 5.65 17.74 39.95
From Table 14, we observe that DELETE is effective in unlearning writer ID information, achieving only 0.40% classification accuracy on the forget set when applied to the base model $M$, and dropping further to 0.00% when applied to the pruned model $M^{*}$. At the same time, it maintains high accuracy on the retain set, demonstrating strong task retention. Although both CER and WER initially degrade for both the forget and retain sets, the performance significantly improves after the fine-tuning stage, showing the potential of DELETE in handwriting recognition unlearning scenarios. Furthermore, the MIA success rate on the forget set is 46%, which is close to the random baseline of 50%, indicating successful unlearning from a privacy perspective.
4.6.5 Writer-ID Confusion (WIC) Experiments
Finally, we conduct experiments using our proposed Writer-ID Confusion (WIC) method. To determine the optimal value for the confusion loss weight $\lambda$, we perform a hyperparameter search over 2,000 training iterations, as shown in Table 15. The results show that varying $\lambda$ from 0.2 to 0.8 yields generally stable performance across the forget, retain, and test sets. Among the results, $\lambda=0.4$ achieves the best balance between effective forgetting and accuracy preservation, and is therefore used in all subsequent WIC experiments.
Table 15: Hyper-parameter search for our method WIC on pruned models.
$\lambda$ Iter. Sparsity Forget Set Retain Set Test Set
ACC CER WER ACC CER WER CER WER
0.2 2,000 7.47% 0.10 1.12 1.54 99.82 2.19 5.71 15.37 39.63
0.4 2,000 7.47% 0.00 0.81 1.00 99.82 1.79 5.09 15.57 39.80
0.6 2,000 7.46% 0.00 0.59 1.35 99.69 1.90 6.09 15.81 40.83
0.8 2,000 7.46% 0.00 0.42 1.30 99.54 1.80 6.25 15.35 40.13
4.6.6 Final Comparison with State-of-the-Art Methods
Since our proposed WIC method does not require an additional fine-tuning stage, we directly report the prune-WIC results in Table 16. As shown, even with just 2,000 iterations, WIC achieves state-of-the-art performance. When extended to 10,000 iterations, it delivers results that are comparable to or even better than existing methods. To assess the statistical significance of the observed performance differences, we perform a Multivariate Analysis of Variance (MANOVA) using the joint metrics of ACC, CER, and WER, and confirm that the improvements are statistically significant. Furthermore, the MIA success rates on the forget set are 49% and 48% after 2,000 and 10,000 iterations, respectively. Both are close to the random guess baseline of 50%, indicating effective unlearning from a privacy standpoint.
Table 16: Final comparison with state-of-the-art methods on the IAM datasets.
Method Iter. Sparsity Forget Set Retain Set Test Set
ACC CER WER ACC CER WER CER WER
Baseline $M$ 0 0% 100.00 1.80 1.89 100.00 1.29 1.75 13.23 34.90
Pruned $M^{*}$ 0 17.45% 20.03 58.29 84.16 35.50 60.44 85.26 66.81 88.32
Prune-RL 10,000 7.41% 0.00 0.28 1.05 99.95 2.46 3.27 14.98 37.61
Prune-Fisher-FT 10,000 7.39% 8.87 5.21 9.77 99.99 2.52 3.82 15.16 38.14
Prune-Amnesiac-FT 10,000 7.40% 7.47 4.88 8.87 99.99 3.92 4.87 16.53 38.77
Prune-DELETE-FT 10,000 7.43% 0.00 7.01 11.16 99.99 4.70 5.65 17.74 39.95
Prune-WIC (Ours) 2,000 7.47% 0.00 0.81 1.00 99.82 1.79 5.09 15.57 39.80
Prune-WIC (Ours) 10,000 7.42% 0.00 0.97 1.35 99.96 1.75 3.05 14.83 37.38
4.6.7 Final Comparison on the CVL datasets
The CVL datasets comprise a training set of 27 writers and a test set of 283 writers. To evaluate the robustness of our proposed method, we merge all the writers, resulting in a total of 310 writers and 99,904 word-level images. This combined dataset is used for unlearning comparison. We limit the unlearning process to a maximum of 4,000 iterations. The final performance comparison is presented in Table 17.
Our proposed method, Prune-WIC, achieves a compelling balance between effective forgetting and performance retention. On the Forget Set, Prune-WIC attains the lowest writer identification accuracy (ACC) of 0.00 and a WER of 4.68%, outperforming all baseline methods. In particular, it shows significant improvement compared to Prune-Fisher-FT and Prune-DELETE-FT, which suffer from higher CER and WER values. On the Retain Set, Prune-WIC maintains strong retention performance with an accuracy of 98.36%, a CER of 5.70%, and a WER of 7.16%, exhibiting only marginal degradation compared to the original Baseline model.
To assess the statistical significance of the observed performance differences, we perform a Multivariate Analysis of Variance (MANOVA) using the joint metrics of ACC, CER, and WER, and confirm that the improvements are statistically significant.
In terms of computational efficiency, Prune-WIC completes the unlearning process in 4,069 seconds while using 16.4 GB of GPU memory, demonstrating a favorable trade-off between effectiveness and resource consumption. These results confirm the effectiveness and practicality of our approach for unlearning in handwriting recognition scenarios.
Table 17: Final comparison with state-of-the-art methods on the CVL datasets.
Method Iter. Sparsity Forget Set Retain Set Time Peak Mem.
ACC CER WER ACC CER WER (second) GPU (GB)
Baseline $M$ 0 0% 100.00 4.91 5.22 100.00 6.47 5.70 - -
Pruned $M^{*}$ 0 17.44% 5.29 65.40 86.47 14.33 68.70 88.00 3432 13.8
Prune-RL 4,000 7.40% 0.25 3.81 6.55 92.04 6.93 10.35 6362 16.4
Prune-Fisher-FT 4,000 7.30% 0.00 42.72 72.66 33.17 42.26 72.17 1025 16.4
Prune-Amnesiac-FT 4,000 7.40% 9.17 3.68 8.00 96.49 4.21 7.36 1029 16.4
Prune-DELETE-FT 4,000 7.40% 0.00 7.34 10.62 97.47 8.21 9.24 1740 17.8
Prune-WIC (Ours) 4,000 7.39% 0.00 4.68 6.08 98.36 5.70 7.16 4069 16.4
5 Conclusion
In this work, we propose a novel two-stage machine unlearning framework for transformer-based handwriting text recognition (HTR) systems, focusing on removing user-identifiable information while preserving model utility. The first stage introduces a targeted neural pruning strategy that analyzes neuron importance based on activation ratios between the forget and retain sets. This method not only enables effective removal of sensitive information but also provides practical insights into where and how to prune the model. Importantly, this pruning strategy can serve as a general guideline for future unlearning tasks, eliminating the need to repeatedly re-run expensive experiments to determine which blocks to prune or at what probability.
In the second stage, we introduce Writer-ID Confusion (WIC), our key contribution to machine unlearning. WIC enforces a uniform distribution over writer predictions for the forget set while retaining standard training on the retain set. This lightweight, plug-and-play method achieves competitive or superior performance compared to state-of-the-art unlearning techniques, all without requiring additional fine-tuning or excessive computation. WIC proves to be a practical and effective solution for real-world privacy-preserving HTR applications.
While our work primarily focuses on writer identity, it does not yet address other sources of information leakage such as rare lexical patterns or unique stylistic traits. Future research should explore these aspects using techniques like stylometry or lexical distribution analysis to broaden privacy coverage.
Additionally, this framework opens several promising research directions. Future work could explore layout-aware machine unlearning for selectively forgetting content in documents with complex spatial structures, or multilingual unlearning to address language-specific forgetting and cross-linguistic effects. Applying these approaches to downstream tasks such as document visual question answering (DocVQA) [38] could enhance privacy by enabling systems to forget entire documents from retrievable memory. Integrating the proposed prune-unlearn mechanisms into retrieval-based DocVQA systems [39] would allow models to be updated upon user requests, preventing access to specific documents and ensuring stronger end-to-end privacy guarantees.
@@ Acknowledgements
This work has been supported by the Beatriu de Pinós del Departament de Recerca i Universitats de la Generalitat de Catalunya (2022 BP 00256), the predoctoral program AGAUR-FI ajuts (2024 FI-3 00065) Joan Oró, which is backed by the Secretariat of Universities and Research of the Department of Research and Universities of the Generalitat of Catalonia, as well as the European Social Plus Fund, European Lighthouse on Safe and Secure AI (ELSA) from the European Union’s Horizon Europe programme under grant agreement No 101070617, Ramon y Cajal research fellowship RYC2020-030777-I / AEI / 10.13039/501100011033. This work has also been supported by Riksbankens Jubileumsfond, grant M24-0028 (Echoes of History: Analysis and Decipherment of Historical Writings, DESCRYPT), the Spanish projects CNS2022-135947 (DOLORES), PID2021-126808OB-I00 (GRAIL) and PID2024-157778OB-I00 (SUKIDI), the Consolidated Research Group 2021 SGR 01559 from the Research and University Department of the Catalan Government, and PID2023-146426NB-100 funded by MCIU/AEI/10.13039/501100011033 and FSE+.